OpenStudy (anonymous):

sin 3x cos x - cos 3x sin x=.5

OpenStudy (phi):

do you know that sin(A-B)= sin(A)cos(B) - cos(A)sin(B) if you let A= 3x and B= x, can you finish?

OpenStudy (anonymous):

that's were i got stuck i dont know what to do from there

OpenStudy (phi):

OK. start with sin(A)cos(B) - cos(A)sin(B) and match it to your problem sin 3x cos x - cos 3x sin x what is A and what is B?

OpenStudy (anonymous):

3x is A ad x is B

OpenStudy (phi):

so we replace sin(A)cos(B) - cos(A)sin(B) with sin(A-B), but replace A with 3x and B with x

OpenStudy (anonymous):

Ok. but what do you do now

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

i got sin 2x

OpenStudy (phi):

OK, so you have simplified the problem from sin 3x cos x - cos 3x sin x=.5 to sin(2x)= 0.5 Do you know the sin of what angle is 1/2 ?

OpenStudy (phi):

You could use a calculator, but generally people memorize a few key angles.

OpenStudy (anonymous):

pi/6 and 5pi/6

OpenStudy (phi):

so that means if we start with sin(2x)= 1/2 then 2x = pi/6 or 2x= 5pi/6 can you solve for x?

OpenStudy (anonymous):

pi/3 and 5pi/3

OpenStudy (phi):

to solve for x you multiply by 1/2 on both sides: \[\frac{1}{\cancel{2}}\cancel{2}x=\frac{1}{2}\frac{\pi}{6}\] when you multiply the fractions on the right hand side, multiply top*top and bottom*bottom

OpenStudy (phi):

you can check your answer by copy and paste this into the google search window sin(pi/4)*cos(pi/12)- cos( pi/4 )*sin(pi/12) =